How many grams of H3PO4 are produced when 10.0 moles of water react with an excess of P4O10

1
Expert's answer
2021-01-11T03:50:51-0500

P4O10 (s) + 6 H2O (l) "\to" 4 H3PO4 (aq)


10 mole H2O("\frac{4mole H_3PO_4}{6moleH_2O}" ) =

40/6 mole of H3PO4

Mass of H3PO4 = "\frac{40}{6}" ×98 grams = 653.33 g